# AdS3 black holes with primary Proca hair from regularized Gauss-Bonnet coupling

### v3

@@ Abstract
We construct a consistent three-dimensional Einstein-Gauss-Bonnet theory as a vector-tensor theory within the generalized Proca class by employing a regularization procedure based on the Weyl geometry, which was introduced recently in [arXiv:2504.13084]. We then obtain an asymptotically AdS3, static, and circularly symmetric black hole solution with primary Proca hair. Afterwards, we investigate the effect of the scalar-tensor Gauss-Bonnet coupling constructed previously by different regularization schemes. We further generalize these solutions by incorporating an electric charge. As special cases, we find a regular black hole solution in addition to charged and uncharged stealth BTZ black hole solutions.
@@ 1 Introduction
One possible workaround for higher-derivatives in a Lagrangian is to have a trivial degeneracy, which is realized when the Lagrangian in question differ from another Lagrangian with at most first derivatives, only by a total derivative. General relativity is an example scenario where the Einstein-Hilbert action can be written in the following form
$\sqrt{-g}R=\sqrt{-g}g^{ab}\left[\Gamma^{c}_{\,ad}\Gamma^{d}_{\,bc}-\Gamma^{c}_{\,dc}\Gamma^{d}_{\,ab}\right]+\partial_{a}\left[\sqrt{-g}\left(g^{bc}\Gamma^{d}_{\,bc}-g^{ab}\Gamma^{c}_{\,bc}\right)\right],$ (1.1)
which is the Dirac’s $\Gamma\Gamma$ formulation. Although the covariance is lost, one clearly understands why the resulting field equations, the Einstein equations, are second-order differential equations.
@@
$\mathcal{L}_{n}=\frac{1}{2^{n}}\delta_{c_{1}d_{1}\ldots c_{n}d_{n}}^{a_{1}b_{1}\ldots a_{n}b_{n}}R_{a_{1}b_{1}}^{c_{1}d_{1}}\cdots R_{a_{n}b_{n}}^{c_{n}d_{n}},$ (1.2)
@@
where $n$ is the order of the Lovelock Lagrangian and $n=1$ corresponds to the Einstein-Hilbert case. The somewhat dissapointing part of the story is that the $n$-th order Lovelock Lagrangian contributes to the dynamics only when [^1] For $d<2n$, $\mathcal{L}_{n}=0$ due to full anti-symmetry of the generalized Kronecker delta. $d=2n$ is the critical dimension where $\mathcal{L}_{n}$ is a topological invariant. $d>2n$, which leaves the Einstein-Hilbert action as the only possibility in $d=4$. In order to get a non-trivial contribution to field equations from the leading correction, i.e., the Gauss-Bonnet invariant which is given by
$\mathcal{L}_{n=2}=\mathcal{G}=R^{2}-4R_{ab}R^{ab}+R_{abcd}R^{abcd},$ (1.3)
one should work in $d\geq 5$. Although it does not provide a generalization of general relativity in $d=4$, Einstein-Gauss-Bonnet gravity defined with the Lagrangian $\mathcal{L}=R+\alpha\,\mathcal{G}$ ($\alpha$: constant) provides an excellent opportunity to test any theoretical idea beyond general relativity. Fortunately, scalar-tensor theories and vector-tensor theories that are also free of Ostragadski instabilities, called as Horndeski theories [5, 6] and generalized Proca theories [7] respectively, are available for phenomenological applications in $d=4$.
@@
Recently, Glavan and Lin proposed a procedure to define the Einstein-Gauss-Bonnet theory in $d<5$ [8]. Studying the solutions of the theory in $d$-dimensions by using an appropriate ansatz, one finds that the contribution of the Gauss-Bonnet term to field equations come with a factor of (d3)(d4)(d-3)(d-4). If one scales the coupling as $\alpha\to\frac{\alpha}{d-p}$ ($p=3,4$) and then set $d=p$, it is possible to get novel solutions in $d=3,4$. While it works for constant curvature spacetimes, cosmological spacetimes and spherically symmetric black holes, it turns out that the field equations for a general metric contains a part which is always higher-dimensional [9]. Therefore, one does not get a consistent Einstein-Gauss-Bonnet theory in $d=3,4$ through this procedure.
@@
Although this initial attempt failed, well-defined regularization schemes were discovered [10, 11, 12, 13], which was also applied to the third-order Lovelock Lagrangian [14]. In order to obtain a consistent lower-dimensional theory, one can either use the difference of the Gauss-Bonnet actions corresponding to two conformally-related metrics or perform a Kaluza-Klein reduction with an internal space conformally related to a maximally symmetric space. By again scaling the coupling and then fixing the dimension at the end, not surprisingly, one finds a scalar-tensor theory withing the Horndeski class. For details of the properties of the resulting theories, the reader is referred to [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. The regularized Lagrangian at the general $n$-th order was obtained in [27]. By considering an infinite tower of corrections to the Einstein-Hilbert action in $d=4$, which leads to non-local Horndeski theories, it is possible to obtain cosmological models with no big bang singularity, and also regular black hole solutions. Thermodynamics of black hole solutions of these theories was studied in [28]. Regular black hole solutions in $d=3$ were given in [29].
@@
The sign in front of the square-root can be fixed to be minus by demanding a well-defined $\alpha\to 0$ limit. Remarkably, the metric function $f$ takes the same form as one gets from the procedure of Glavan and Lin. However, it is now realized as a solution of a particular 3d Horndeski theory and a non-trivial scalar profile supports the solution. Note that there appears no integration constant in the metric function due to the scalar hair in addition the constant $m$, which is related to the mass of the black hole (see [16] for details) and the geometry is modified only by the coupling constant $\alpha$. Therefore, it is a black hole with secondary hair.
@@
An arguably more interesting scenario would be to have a black hole solution with primary hair, where an additional integration constant emerges in the metric function. In such a case, considerable deviations from the solutions of general relativity can be obtained even when the coupling constant is small. However, such solutions are quite rare in the literature. Very recently, by introducing another regularization scheme for the Gauss-Bonnet term within the framework of Weyl geometry, a 4d vector-tensor theory of the generalized Proca class was obtained [30]. This theory admits asymptotically flat, static, and spherically symmetric black hole solutions with primary hair. For the study of particle motion, shadows, and grey-body factors one may refer to [31].
@@
In this work, motivated by various interesting properties of the scalar-tensor 3d Einstein-Gauss-Bonnet theory presented in the literature [15, 16, 17, 21, 25], we construct the vector-tensor 3d Einstein-Gauss-Bonnet theory by using the regularization of [30] and give asymptotically AdS3, static, and circularly symmetric black hole solutions with primary hair. The outline of this paper is as follows: In Section 2, we review the procedure of [30] for the regularization of the Gauss-Bonnet term and present the resulting vector-tensor Lagrangian in $d=3$. In Section 3, we derive a black hole solution with primary hair by using this Lagrangian as a correction to the Einstein-Hilbert action. Then, we generalize this solution to a case where the effect of the scalar-tensor Gauss-Bonnet term in (1.4) is taken into account. We also find the electrically charged version of the most general solution and we end our paper with a summary of our results followed by some directions for future research in Section 4.
@@
The logic behind the regularization of [30] is as follows: The Gauss-Bonnet invariant does not contribute to the dynamics when $d<5$ independent of the connection that is used. Therefore, when we calculate the difference between the Gauss-Bonnet term with the Weyl connection and with the Levi-Civita connection, the difference cannot contribute to the dynamics either, and therefore can only contain terms with (d3)(d4)(d-3)(d-4) up to boundary terms. Then, a nontrivial Lagrangian contributing to field equations in lower-dimensions can be obtained by scaling the coupling and then fixing the dimension at the end as before.
For a torsionless Weyl connection $\widetilde{\Gamma}^{\alpha}_{\ \mu\nu}$ with the corresponding covariant derivative $\widetilde{\nabla}_{\mu}$, the covariant derivative of the metric does not vanish but take the following form
@@
In [30], the Gauss-Bonnet term calculated with the Weyl connection was written in the form which allows a regularization as follows
@@ 3 Black hole solutions with primary hair
$L_{\text{red}}=$ $-\frac{r\beta w_{0}^{4}}{f^{2}N^{3}}+\frac{2r\beta w_{0}^{2}w_{1}N^{\prime}}{N^{2}}+2\beta f^{2}w_{1}^{2}(1+rw_{1})N^{\prime}+\frac{\beta w_{0}^{2}\left[(1+2rw_{1})f^{\prime}+2f(w_{1}+rw_{1}^{2}+rw_{1}^{\prime})\right]}{fN}$
@@
The equations of motion can be found by extremizing the reduced action $I_{\text{red}}$. Unlike what we have for the scalar-tensor version of the 3d Einstein-Gauss-Bonnet theory, it is not possible to integrate the equations following from the variation of the action directly. However, as in the 4d case [30], a close inspection of the equations show that we have the following symmetry
@@
where $c$ is another integration constant which does not modify the geometry. The norm of the vector field is constant on-shell, i.e. $W^{2}=-2c$. Using the symmetry (3.4) with $\kappa=-c$, one gets the same expressions for the components of the vector field with $c=0$. This explains why this constant has no effect on the geometry.
@@
is the effective cosmological constant. In order to have an asymptotically AdS3 solution ($\Lambda<0$), the following constaints on the parameters should be satisfied
@@
As we see, the bare cosmological constant $\Lambda_{0}$ in the action (2.8) should always be negative and the coupling constant $\beta$ should be positive but has an upper bound depending on the bare cosmological constant. In Figure 1, the effect of the hair parameter $q$ is shown for a particular choice of $(m,\Lambda_{0},\beta)$ satisfying the constaints. One always has a black hole with a single horizon and increasing the value of $q$ shifts the location of the event horizon towards the larger values of $r$. As apperant from (3.10), while $w_{0}^{2}$ is finite at the event horizon, $w_{1}$ diverges.
@@ 3.2 Addition of an electric charge
for $\alpha\neq\beta$. The vector components in terms of the metric function $f$ remains the same as in (3.10).
@@ 4 Summary and outlook
In this paper, we constructed the 3d Einstein-Gauss-Bonnet theory as a vector-tensor theory within the generalized Proca class by using the regularization procedure introduced recently in [30]. We studied the static and circularly symmetric black hole solutions of this theory. As showed in [30] for 4d, unlike the solutions of the scalar-tensor theories obtained as lower-dimensional limits by different regularization schemes, the resulting equations of motion for the unknown functions can only be solved with the help of the Noether charge associated with a global symmetry of the reduced action. We found an AdS3 black hole with primary hair, which has only a few examples in the literature [40, 41]. We then generalized it by introducing the scalar-tensor Gauss-Bonnet invariant into the action. When the coupling constants of the scalar-tensor and vector-tensor invariants are equal, we obtained a regular black hole solution. Additionally, we presented the electrically charged version and as a by-product, we observed that stealth BTZ and charged BTZ black hole solutions arise for an appropriate choice of the hair parameter.
Our work opens up quite a few possibilities for future work. First of all, in order to understand the physical significance of the hair parameter $q$, thermodynamics of these solutions should be studied. As in the case of the scalar-tensor 3d Einstein-Gauss-Bonnet theory [17], rotating black hole solutions can be obtained by boosting static solutions [42, 43]. The scalar-tensor 3d Einstein-Gauss-Bonnet theory together with its cubic extension constructed in [14] exhibit remarkable holographic properties. They were shown to admit a holographic c-theorem [21] and it is possible to give a microscopic derivation of the entropy of their static black hole solutions [44]. Therefore, it would be interesting to construct the cubic extension of the vector-tensor version and check whether they have similar properties. Very recently, an alternative regularization mechanism has been proposed[45], where the difference between Gauss-Bonnet invariants with two different Weyl vectors is used. The resulting bi-Proca theory gives rise to regular black hole solutions. A similar behavior is expected in 3d.

### v4

# AdS3 black holes with primary Proca hair from a regularized Gauss-Bonnet coupling

@@ Abstract
We construct a consistent three-dimensional Einstein-Gauss-Bonnet theory as a vector-tensor theory within the generalized Proca class by employing a regularization procedure based on the Weyl geometry, which was introduced recently in [arXiv:2504.13084]. We then obtain an asymptotically AdS3, static, and circularly symmetric black hole solution with primary Proca hair. Afterward, we investigate the effect of the scalar-tensor Gauss-Bonnet coupling constructed previously by different regularization schemes. We further generalize these solutions by incorporating an electric charge. As special cases, we find a regular black hole solution in addition to charged and uncharged stealth BTZ black hole solutions.
@@ 1 Introduction
One possible workaround for higher-derivatives in a Lagrangian is to have a trivial degeneracy, which is realized when the Lagrangian in question differs from another Lagrangian with at most first derivatives, only by a total derivative. General relativity (GR) is an example scenario where the Einstein-Hilbert action can be written in the following form
$\sqrt{-g}R=\sqrt{-g}g^{\mu\nu}\left[\Gamma^{\rho}_{\,\mu\sigma}\Gamma^{\sigma}_{\,\nu\rho}-\Gamma^{\rho}_{\,\sigma\rho}\Gamma^{\sigma}_{\,\mu\nu}\right]+\partial_{\mu}\left[\sqrt{-g}\left(g^{\nu\rho}\Gamma^{\mu}_{\,\nu\rho}-g^{\mu\nu}\Gamma^{\rho}_{\,\nu\rho}\right)\right],$ (1.1)
which is the so-called $\Gamma\Gamma$ form of the Lagrangian of GR that he used in his Hamiltonian formulation [2]. Although the covariance is lost, one clearly understands why the resulting field equations, the Einstein equations, are second-order differential equations.
@@
$\mathcal{L}_{n}=\frac{1}{2^{n}}\delta_{\rho_{1}\sigma_{1}\ldots\rho_{n}\sigma_{n}}^{\mu_{1}\nu_{1}\ldots\mu_{n}\nu_{n}}R_{\ \ \ \ \mu_{1}\nu_{1}}^{\rho_{1}\sigma_{1}}\cdots R_{\ \ \ \ \mu_{n}\nu_{n}}^{\rho_{n}\sigma_{n}},$ (1.2)
@@
where $n$ is the order of the Lovelock Lagrangian and $n=1$ corresponds to the Einstein-Hilbert case. The somewhat disappointing part of the story is that the $n$-th order Lovelock Lagrangian contributes to the dynamics only when [^1] For $d<2n$, $\mathcal{L}_{n}=0$ due to full anti-symmetry of the generalized Kronecker delta. $d=2n$ is the critical dimension where $\mathcal{L}_{n}$ is a topological invariant. $d>2n$, which leaves the Einstein-Hilbert action as the only possibility in $d=4$. In order to get a non-trivial contribution to field equations from the leading correction, i.e., the Gauss-Bonnet invariant which is given by
$\mathcal{L}_{n=2}=\mathcal{G}=R^{2}-4R_{\mu\nu}R^{\mu\nu}+R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma},$ (1.3)
one should work in $d\geq 5$. Although it does not provide a generalization of GR in $d=4$, Einstein-Gauss-Bonnet gravity defined with the Lagrangian $\mathcal{L}=R+\alpha\,\mathcal{G}$, where $\alpha$ is a constant, provides an excellent opportunity to test quite a few theoretical ideas beyond GR, especially in the context of AdS/CFT correspondence [6, 7, 8]. Most notably, the holographic studies of the theory in $d\geq 5$ have yielded important results about the micro-causality and the hydrodynamic properties of the dual conformal field theories defined in (d1)(d-1)-dimensions [9, 10, 11, 12, 13, 14, 15].
If one does not insist on working with a pure gravity theory, one can construct scalar-tensor and vector-tensor theories in $d<5$ that are also free of Ostragadski instabilities, which are called Horndeski theories [16, 17] and generalized Proca theories [18] respectively. Having second-order field equations, all these theories are natural candidates for admitting exact black hole solutions. As will be explained soon, through different regularization procedures, one can obtain examples of Horndeski or generalized Proca theories starting from Lovelock Lagrangians in $d\geq 5$.
After it was realized that GR with a negative cosmological constant admits an AdS3 black hole solution, the Banados-Teitelboim-Zanelli (BTZ) black hole [19, 20], gravity in three dimensions has become an active field of research since it allows us to study black hole physics in relatively simple toy models. In particular, we can probe the properties of quantum gravity theories defined in AdS3 spacetime through the AdS3/CFT2 correspondence (see [21] for a review) and the infinite number of symmetries present in a 2d CFT [22] is a particularly powerful tool in this regard. For example, we can give a microscopic derivation of the semi-classical entropy of the BTZ black hole by using the Cardy formula for the asymptotic growth of the number of states in a 2d CFT [23]. Checking whether this can be achieved beyond GR, with solutions different than the BTZ black hole if possible, is a non-trivial test for the generality of the AdS3/CFT2 correspondence. Therefore, it is of importance to obtain AdS3 black hole solutions in different setups including various matter couplings and higher-curvature extensions.
@@
Recently, Glavan and Lin proposed a procedure to define the Einstein-Gauss-Bonnet theory in $d<5$ [24]. Studying the solutions of the theory in $d$ dimensions by using an appropriate ansatz, one finds that the contribution of the Gauss-Bonnet term to field equations comes with a factor of (d3)(d4)(d-3)(d-4). If one scales the coupling constant as $\alpha\to\frac{\alpha}{d-p}$ ($p=3,4$) and then sets $d=p$, it is possible to get novel solutions in $d=3,4$. While it works for constant curvature spacetimes, cosmological spacetimes and spherically symmetric black holes, it turns out that the field equations for a general metric contain a part which is always higher-dimensional [25]. Therefore, one does not get a consistent Einstein-Gauss-Bonnet theory in $d=3,4$ through this procedure.
@@
Although this initial attempt failed, well-defined regularization schemes were discovered [26, 27, 28, 29], which was also applied to the third-order Lovelock Lagrangian [30]. In order to obtain a consistent lower-dimensional theory, one can either use the difference of the Gauss-Bonnet actions corresponding to two conformally related metrics or perform a Kaluza-Klein reduction with an internal space conformally related to a maximally symmetric space. By again scaling the coupling and then fixing the dimension at the end, not surprisingly, one finds a scalar-tensor theory withing the Horndeski class. For details of the properties of the resulting theories, the reader is referred to [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. The regularized Lagrangian at the general $n$-th order was obtained in [43]. By considering an infinite tower of corrections to the Einstein-Hilbert action in $d=4$, which leads to non-local Horndeski theories, it is possible to obtain cosmological models with no big bang singularity, and also regular black hole solutions. Thermodynamics of black hole solutions of these theories was studied in [44]. Regular black hole solutions in $d=3$ were given in [45].
@@
The sign in front of the square root can be fixed to be minus by demanding a well-defined $\alpha\to 0$ limit. Remarkably, the metric function $f$ takes the same form as one gets from the procedure of Glavan and Lin. However, it is now realized as a solution of a particular 3d Horndeski theory and a non-trivial scalar profile supports the solution. Note that there appears no integration constant in the metric function due to the scalar hair in addition the constant $m$, which is related to the mass of the black hole (see [32] for details) and the geometry is modified only by the coupling constant $\alpha$. Therefore, it is a black hole with secondary hair.
For black hole solutions with a scalar hair, a microscopic derivation of the semi-classical entropy is still possible. However, the Cardy formula must first be written in terms of the ground state energy instead of the central charge of the theory. If the ground state of the dual 2d CFT is identified with the soliton solution, which is obtained from the static black hole solution by a double Wick rotation, the semi-classical entropy can be reproduced [46, 47, 48]. This idea was successfully applied to the solution (1.7) of the regularized Einstein-Gauss-Bonnet theory in $d=3$. The semi-classical entropy obeys the area law ($S=\frac{A}{4G}$ where $A$ is the area of the event horizon) and it was shown to follow from the alternative Cardy formula of [46, 47, 48] in [49]. 3d Cubic Lovelock gravity obtained in [30] is a more interesting setup. In addition to a black hole solution whose entropy obeys the area law, the theory also admits the BTZ black hole as a solution with a deviation from the area law. In both cases, the alternative Cardy formula works if the corresponding soliton solutions are identified with distinct ground states [49]. This shows that these two black hole solutions belong to disconnected sectors of the theory and each sector has its own ground state.
@@
An arguably more interesting scenario would be to have a black hole solution with primary hair, where an additional integration constant emerges in the metric function. In such a case, considerable deviations from the solutions of GR can be obtained even when the coupling constant is small. However, such solutions are quite rare in the literature. Very recently, by introducing another regularization scheme for the Gauss-Bonnet term within the framework of Weyl geometry, a 4d vector-tensor theory of the generalized Proca class was obtained [50]. This theory admits asymptotically flat, static, and spherically symmetric black hole solutions with primary hair. For the study of particle motion, shadows, and grey-body factors one may refer to [51].
In this work, motivated by the success of the scalar-tensor 3d Einstein-Gauss-Bonnet theory (and its cubic extension) in testing the AdS3/CFT2 correspondence in a highly non-trivial setup, we construct the vector-tensor 3d Einstein-Gauss-Bonnet theory by using the regularization of [50]. Studying static and circularly symmetric black hole solutions, we find asymptotically AdS3 solutions with primary hair analogous to those given in [50]. With these solutions, it becomes possible to check the validity of the prescription of [46, 47, 48] for black holes with primary hair, which we leave as future work.
@@
The outline of this paper is as follows: In Section 2, we review the procedure of [50] for the regularization of the Gauss-Bonnet term and present the resulting vector-tensor Lagrangian in $d=3$. In Section 3, we derive a black hole solution with primary hair by using this Lagrangian as a correction to the Einstein-Hilbert action. Then, we generalize this solution to a case where the effect of the scalar-tensor Gauss-Bonnet term in (1.4) is taken into account. We also find the electrically charged version of the most general solution and we end our paper with a summary of our results followed by some directions for future research in Section 4.
@@
The logic behind the regularization of [50] is as follows: The Gauss-Bonnet invariant does not contribute to the dynamics when $d<5$ independent of the connection that is used. Therefore, when we calculate the difference between the Gauss-Bonnet term with the Weyl connection and with the Levi-Civita connection, the difference cannot contribute to the dynamics either, and therefore can only contain terms with (d3)(d4)(d-3)(d-4) up to boundary terms. Then, a nontrivial Lagrangian contributing to field equations in lower dimensions can be obtained by scaling the coupling and then fixing the dimension at the end as before.
For a torsionless Weyl connection $\widetilde{\Gamma}^{\alpha}_{\ \mu\nu}$ with the corresponding covariant derivative $\widetilde{\nabla}_{\mu}$, the covariant derivative of the metric does not vanish but takes the following form
@@
In [50], the Gauss-Bonnet term calculated with the Weyl connection was written in the form that allows a regularization as follows
@@ 3 Black hole solutions with primary hair
$L_{\text{red}}=$ $-\frac{\beta rw_{0}^{4}}{f^{2}N^{3}}+\frac{2\beta rw_{0}^{2}w_{1}N^{\prime}}{N^{2}}+2\beta f^{2}w_{1}^{2}(1+rw_{1})N^{\prime}+\frac{\beta w_{0}^{2}\left[(1+2rw_{1})f^{\prime}+2f(w_{1}+rw_{1}^{2}+rw_{1}^{\prime})\right]}{fN}$
@@
The equations of motion can be found by extremizing the reduced action $I_{\text{red}}$. Unlike what we have for the scalar-tensor version of the 3d Einstein-Gauss-Bonnet theory, it is not possible to integrate the equations following from the variation of the action directly. However, as in the 4d case [50], a close inspection of the equations shows that we have the following symmetry
@@
where $c$ is another integration constant that does not modify the geometry. The norm of the vector field is constant on-shell, i.e., $W^{2}=-2c$. Using the symmetry (3.4) with $\kappa=-c$, one gets the same expressions for the components of the vector field with $c=0$. This explains why this constant has no effect on the geometry.
@@
is the effective cosmological constant. In order to have an asymptotically AdS3 solution ($\Lambda<0$), the following constraints on the parameters should be satisfied
@@
As we see, the bare cosmological constant $\Lambda_{0}$ in the action (2.8) should always be negative and the coupling constant $\beta$ should be positive but has an upper bound depending on the bare cosmological constant. In Figure 1, the effect of the hair parameter $q$ is shown for a particular choice of $(m,\Lambda_{0},\beta)$ satisfying the constraints. One always has a black hole with a single horizon and increasing the value of $q$ shifts the location of the event horizon toward the larger values of $r$. As apparant from (3.10), while $w_{0}^{2}$ is finite at the event horizon, $w_{1}$ diverges.
@@ 3.2 Addition of an electric charge
for $\alpha\neq\beta$. The vector components in terms of the metric function $f$ remain the same as in (3.10).
@@ 4 Summary and outlook
In this paper, we constructed the 3d Einstein-Gauss-Bonnet theory as a vector-tensor theory within the generalized Proca class by using the regularization procedure introduced recently in [50]. We studied the static and circularly symmetric black hole solutions of this theory. As shown in [50] for 4d, unlike the solutions of the scalar-tensor theories obtained as lower-dimensional limits by different regularization schemes, the resulting equations of motion for the unknown functions can only be solved with the help of the Noether charge associated with a global symmetry of the reduced action. We found an AdS3 black hole with primary hair, which has only a few examples in the literature [59, 60]. We then generalized it by introducing the scalar-tensor Gauss-Bonnet invariant into the action. When the coupling constants of the scalar-tensor and vector-tensor invariants are equal, we obtained a regular black hole solution. Additionally, we presented the electrically charged version and as a by-product, we observed that stealth BTZ and charged BTZ black hole solutions arise for an appropriate choice of the hair parameter.
Our work opens up quite a few possibilities for future work. First of all, in order to understand the physical significance of the hair parameter $q$, thermodynamics of these solutions should be studied. As in the case of the scalar-tensor 3d Einstein-Gauss-Bonnet theory [33], rotating black hole solutions can be obtained by boosting static solutions [61, 62]. The scalar-tensor 3d Einstein-Gauss-Bonnet theory together with its cubic extension constructed in [30] exhibit remarkable holographic properties. As mentioned in the introduction, it is possible to give a microscopic derivation of the entropy of their static black hole solutions [49]. Additionally, for these theories, one can construct a monotonically increasing c-function whose value at the UV fixed point matches with the Weyl anomaly coefficient [37]. This means that they admit a holographic c-theorem [63, 64, 65], which mimics the Zamolodchikov’s c-theorem for 2d QFTs. Theories admitting a holographic c-theorem are quite rare and the only other 3d examples are new massive gravity [66, 67], its certain higher-curvature [56] and Born-Infeld type [68, 69, 70] extensions. A natural question is whether these properties are shared by the 3d vector-tensor Einstein-Gauss-Bonnet theory that we constructed here and its cubic extension. Since the black hole solutions presented here possess primary hair, the corresponding ground states used in the microscopic derivation of the semi-classical entropy might exhibit particularly interesting properties not observed before. Very recently, an alternative regularization mechanism has been proposed[71], where the difference between Gauss-Bonnet invariants with two different Weyl vectors is used. The resulting bi-Proca theory gives rise to regular black hole solutions. A similar behavior is expected in 3d.